OpenStudy (ashy98):

Can someone help me with 2 questions?

OpenStudy (ashy98):

Simplify open parentheses x to the 2 third power close parentheses to the 4 fifths power

OpenStudy (ashy98):

A.) \[x^\frac{ 2 }{ 5 }\] B.) \[x^\frac{ 2 }{ 15 }\] C.) \[x^\frac{ 8 }{ 15 }\] D.) \[x^\frac{ 22 }{ 15 }\]

OpenStudy (ashy98):

Simplify 64 to the negative 1 third power A.) −4 B.) negative 1 over 4 C.) 1 over 4 D.) 4

Directrix (directrix):

Is this the first problem: